average age of a boy and a girl is 14 yrs. The average age of the same boy, the girl and her twin sister is 16 yrs. What is the boy's age?
Answer by amfagge92(93)

let their ages be a , b and 2b
a+b/2=14....i
a+2b/3=16..ii
solving we have
a+b=28....iii
a+2b=48...iv
a=28-b......v
subtitute a in eqn iv
28-b+2b=48
b=20
subtitute b in enq v
a=28-20
a=8.
the boy's age is 8.